\begin{thebibliography}{35}
\providecommand{\natexlab}[1]{#1}

\bibitem[{Amirkhani et~al.(2023)Amirkhani, AzariJafari, Faridan-Jahromi, Kouhkan, Pourjafari, and Amirak}]{farstail_ds}
Hossein Amirkhani, Mohammad AzariJafari, Soroush Faridan-Jahromi, Zeinab Kouhkan, Zohreh Pourjafari, and Azadeh Amirak. 2023.
\newblock \href {https://doi.org/10.1007/s00500-023-08959-3} {Farstail: a persian natural language inference dataset}.
\newblock \emph{Soft Computing}.

\bibitem[{Ayoubi(2021)}]{pqa_ds}
Mohammad~Yasin Ayoubi, Sajjad \&~Davoodeh. 2021.
\newblock Persianqa: a dataset for persian question answering.
\newblock \url{https://github.com/SajjjadAyobi/PersianQA}.

\bibitem[{Azizi et~al.(2024)Azizi, Chegeni, Mousavi, and Hashemi}]{shiraz_tehran}
Mehrdad Azizi, Reza~Salehi Chegeni, Parisa Mousavi, and Iman Hashemi. 2024.
\newblock [optimizing pre-trained bert-based models for persian language processing].

\bibitem[{Brown et~al.(2020)Brown, Mann, Ryder, Subbiah, Kaplan, Dhariwal, Neelakantan, Shyam, Sastry, Askell, Agarwal, Herbert-Voss, Krueger, Henighan, Child, Ramesh, Ziegler, Wu, Winter, Hesse, Chen, Sigler, Litwin, Gray, Chess, Clark, Berner, McCandlish, Radford, Sutskever, and Amodei}]{NEURIPS2020_1457c0d6}
Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared~D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel Ziegler, Jeffrey Wu, Clemens Winter, Chris Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020.
\newblock \href {https://proceedings.neurips.cc/paper_files/paper/2020/file/1457c0d6bfcb4967418bfb8ac142f64a-Paper.pdf} {Language models are few-shot learners}.
\newblock In \emph{Advances in Neural Information Processing Systems}, volume~33, pages 1877--1901. Curran Associates, Inc.

\bibitem[{Conneau et~al.(2020)Conneau, Khandelwal, Goyal, Chaudhary, Wenzek, Guzmán, Grave, Ott, Zettlemoyer, and Stoyanov}]{conneau_unsupervised_2020}
Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume Wenzek, Francisco Guzmán, Edouard Grave, Myle Ott, Luke Zettlemoyer, and Veselin Stoyanov. 2020.
\newblock \href {https://doi.org/10.48550/arXiv.1911.02116} {Unsupervised {Cross}-lingual {Representation} {Learning} at {Scale}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:1911.02116 [cs].

\bibitem[{Cui et~al.(2021)Cui, Che, Liu, Qin, and Yang}]{cui_pre-training_2021}
Yiming Cui, Wanxiang Che, Ting Liu, Bing Qin, and Ziqing Yang. 2021.
\newblock \href {https://doi.org/10.1109/TASLP.2021.3124365} {Pre-{Training} with {Whole} {Word} {Masking} for {Chinese} {BERT}}.
\newblock \emph{IEEE/ACM Transactions on Audio, Speech, and Language Processing}, 29:3504--3514.
\newblock ArXiv:1906.08101 [cs].

\bibitem[{Dao(2023)}]{dao_flashattention-2_2023}
Tri Dao. 2023.
\newblock \href {https://doi.org/10.48550/arXiv.2307.08691} {{FlashAttention}-2: {Faster} {Attention} with {Better} {Parallelism} and {Work} {Partitioning}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:2307.08691 [cs].

\bibitem[{Darvishi et~al.(2023)Darvishi, Shahbodaghkhan, Abbasiantaeb, and Momtazi}]{pquad_ds}
Kasra Darvishi, Newsha Shahbodaghkhan, Zahra Abbasiantaeb, and Saeedeh Momtazi. 2023.
\newblock \href {https://doi.org/10.1016/j.csl.2023.101486} {Pquad: A persian question answering dataset}.
\newblock \emph{Computer Speech Language}, 80:101486.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and Toutanova}]{devlin_bert_2019}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock \href {https://doi.org/10.18653/v1/N19-1423} {{BERT}: {Pre}-training of {Deep} {Bidirectional} {Transformers} for {Language} {Understanding}}.
\newblock In \emph{Proceedings of the 2019 {Conference} of the {North} {American} {Chapter} of the {Association} for {Computational} {Linguistics}: {Human} {Language} {Technologies}, {Volume} 1 ({Long} and {Short} {Papers})}, pages 4171--4186, Minneapolis, Minnesota. Association for Computational Linguistics.

\bibitem[{Farahani et~al.(2021)Farahani, Gharachorloo, Farahani, and Manthouri}]{parsbert_model}
Mehrdad Farahani, Mohammad Gharachorloo, Marzieh Farahani, and Mohammad Manthouri. 2021.
\newblock \href {https://doi.org/10.1007/s11063-021-10528-4} {Parsbert: Transformer-based model for persian language understanding}.
\newblock \emph{Neural Processing Letters}, 53(6):3831–3847.

\bibitem[{Ghafouri et~al.(2023)Ghafouri, Abbasi, and Naderi}]{ariabert_model}
Arash Ghafouri, Mohammad~Amin Abbasi, and Hassan Naderi. 2023.
\newblock \href {https://doi.org/10.21203/rs.3.rs-3558473/v1} {{AriaBERT}: {A} {Pre}-trained {Persian} {BERT} {Model} for {Natural} {Language} {Understanding}}.
\newblock ISSN: 2693-5015.

\bibitem[{Hosseini et~al.(2018)Hosseini, Ramaki, Maleki, Anvari, and Mirroshandel}]{sp_ds}
Pedram Hosseini, Ali~Ahmadian Ramaki, Hassan Maleki, Mansoureh Anvari, and Seyed~Abolghasem Mirroshandel. 2018.
\newblock Sentipers: A sentiment analysis corpus for persian.
\newblock \emph{arXiv preprint arXiv:1801.07737}.

\bibitem[{Howard and Ruder(2018)}]{howard-ruder-2018-universal}
Jeremy Howard and Sebastian Ruder. 2018.
\newblock \href {https://doi.org/10.18653/v1/P18-1031} {Universal language model fine-tuning for text classification}.
\newblock In \emph{Proceedings of the 56th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, pages 328--339, Melbourne, Australia. Association for Computational Linguistics.

\bibitem[{Khansari and Shamsfard(2021)}]{khansari2021hmblogs}
Hamzeh~Motahari Khansari and Mehrnoush Shamsfard. 2021.
\newblock \href {https://arxiv.org/abs/2111.02362} {Hmblogs: A big general persian corpus}.
\newblock \emph{Preprint}, arXiv:2111.02362.

\bibitem[{Khashabi et~al.(2021)Khashabi, Cohan, Shakeri, Hosseini, Pezeshkpour, Alikhani, Aminnaseri, Bitaab, Brahman, Ghazarian, Gheini, Kabiri, Mahabagdi, Memarrast, Mosallanezhad, Noury, Raji, Rasooli, Sadeghi, Azer, Samghabadi, Shafaei, Sheybani, Tazarv, and Yaghoobzadeh}]{parsinlu_ds}
Daniel Khashabi, Arman Cohan, Siamak Shakeri, Pedram Hosseini, Pouya Pezeshkpour, Malihe Alikhani, Moin Aminnaseri, Marzieh Bitaab, Faeze Brahman, Sarik Ghazarian, Mozhdeh Gheini, Arman Kabiri, Rabeeh~Karimi Mahabagdi, Omid Memarrast, Ahmadreza Mosallanezhad, Erfan Noury, Shahab Raji, Mohammad~Sadegh Rasooli, Sepideh Sadeghi, Erfan~Sadeqi Azer, Niloofar~Safi Samghabadi, Mahsa Shafaei, Saber Sheybani, Ali Tazarv, and Yadollah Yaghoobzadeh. 2021.
\newblock \href {https://doi.org/10.1162/tacl_a_00419} {{P}arsi{NLU}: A suite of language understanding challenges for {P}ersian}.
\newblock \emph{Transactions of the Association for Computational Linguistics}, 9:1147--1162.

\bibitem[{Kudugunta et~al.(2023)Kudugunta, Caswell, Zhang, Garcia, Choquette-Choo, Lee, Xin, Kusupati, Stella, Bapna, and Firat}]{kudugunta2023madlad400}
Sneha Kudugunta, Isaac Caswell, Biao Zhang, Xavier Garcia, Christopher~A. Choquette-Choo, Katherine Lee, Derrick Xin, Aditya Kusupati, Romi Stella, Ankur Bapna, and Orhan Firat. 2023.
\newblock \href {https://arxiv.org/abs/2309.04662} {Madlad-400: A multilingual and document-level large audited dataset}.
\newblock \emph{Preprint}, arXiv:2309.04662.

\bibitem[{Liang et~al.(2023)Liang, Gonen, Mao, Hou, Goyal, Ghazvininejad, Zettlemoyer, and Khabsa}]{liang_xlm-v_2023}
Davis Liang, Hila Gonen, Yuning Mao, Rui Hou, Naman Goyal, Marjan Ghazvininejad, Luke Zettlemoyer, and Madian Khabsa. 2023.
\newblock \href {https://doi.org/10.18653/v1/2023.emnlp-main.813} {{XLM}-{V}: {Overcoming} the {Vocabulary} {Bottleneck} in {Multilingual} {Masked} {Language} {Models}}.
\newblock In \emph{Proceedings of the 2023 {Conference} on {Empirical} {Methods} in {Natural} {Language} {Processing}}, pages 13142--13152, Singapore. Association for Computational Linguistics.

\bibitem[{Liu et~al.(2019)Liu, Ott, Goyal, Du, Joshi, Chen, Levy, Lewis, Zettlemoyer, and Stoyanov}]{liu_roberta_2019}
Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov. 2019.
\newblock \href {https://doi.org/10.48550/arXiv.1907.11692} {{RoBERTa}: {A} {Robustly} {Optimized} {BERT} {Pretraining} {Approach}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:1907.11692 [cs].

\bibitem[{Malmasi et~al.(2022)Malmasi, Fang, Fetahu, Kar, and Rokhlenko}]{multiconer_ds}
Shervin Malmasi, Anjie Fang, Besnik Fetahu, Sudipta Kar, and Oleg Rokhlenko. 2022.
\newblock \href {https://aclanthology.org/2022.coling-1.334} {{M}ulti{C}o{NER}: A large-scale multilingual dataset for complex named entity recognition}.
\newblock In \emph{Proceedings of the 29th International Conference on Computational Linguistics}, pages 3798--3809, Gyeongju, Republic of Korea. International Committee on Computational Linguistics.

\bibitem[{Masumi et~al.(2024)Masumi, Majd, Shamsfard, and Beigy}]{masumi_fabert_2024}
Mostafa Masumi, Seyed~Soroush Majd, Mehrnoush Shamsfard, and Hamid Beigy. 2024.
\newblock \href {https://doi.org/10.48550/arXiv.2402.06617} {{FaBERT}: {Pre}-training {BERT} on {Persian} {Blogs}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:2402.06617 [cs].

\bibitem[{Mirzaee et~al.(2022)Mirzaee, Peymanfard, Moshtaghin, and Zeinali}]{arman_emotion}
Hossein Mirzaee, Javad Peymanfard, Hamid~Habibzadeh Moshtaghin, and Hossein Zeinali. 2022.
\newblock \href {https://doi.org/10.48550/ARXIV.2207.11808} {Armanemo: A persian dataset for text-based emotion detection}.
\newblock \emph{arXiv preprint}.

\bibitem[{Peters et~al.(2018)Peters, Neumann, Iyyer, Gardner, Clark, Lee, and Zettlemoyer}]{peters_deep_2018}
Matthew~E. Peters, Mark Neumann, Mohit Iyyer, Matt Gardner, Christopher Clark, Kenton Lee, and Luke Zettlemoyer. 2018.
\newblock \href {https://doi.org/10.18653/v1/N18-1202} {Deep {Contextualized} {Word} {Representations}}.
\newblock In \emph{Proceedings of the 2018 {Conference} of the {North} {American} {Chapter} of the {Association} for {Computational} {Linguistics}: {Human} {Language} {Technologies}, {Volume} 1 ({Long} {Papers})}, pages 2227--2237, New Orleans, Louisiana. Association for Computational Linguistics.

\bibitem[{Portes et~al.(2023)Portes, Trott, Havens, KING, Venigalla, Nadeem, Sardana, Khudia, and Frankle}]{portes2023mosaicbert}
Jacob Portes, Alexander~R Trott, Sam Havens, DANIEL KING, Abhinav Venigalla, Moin Nadeem, Nikhil Sardana, Daya Khudia, and Jonathan Frankle. 2023.
\newblock \href {https://openreview.net/forum?id=5zipcfLC2Z} {Mosaic{BERT}: A bidirectional encoder optimized for fast pretraining}.
\newblock In \emph{Thirty-seventh Conference on Neural Information Processing Systems}.

\bibitem[{Provilkov et~al.(2020)Provilkov, Emelianenko, and Voita}]{bpe-drop}
Ivan Provilkov, Dmitrii Emelianenko, and Elena Voita. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.170} {{BPE}-dropout: Simple and effective subword regularization}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics}, pages 1882--1892, Online. Association for Computational Linguistics.

\bibitem[{Radford et~al.(2018)Radford, Narasimhan, Salimans, and Sutskever}]{Radford2018ImprovingLU}
Alec Radford, Karthik Narasimhan, Tim Salimans, and Ilya Sutskever. 2018.
\newblock \href {https://openai.com/index/language-unsupervised/} {Improving language understanding by generative pre-training}.

\bibitem[{Rajbhandari et~al.(2020)Rajbhandari, Rasley, Ruwase, and He}]{rajbhandari_zero_2020}
Samyam Rajbhandari, Jeff Rasley, Olatunji Ruwase, and Yuxiong He. 2020.
\newblock \href {https://doi.org/10.48550/arXiv.1910.02054} {{ZeRO}: {Memory} {Optimizations} {Toward} {Training} {Trillion} {Parameter} {Models}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:1910.02054 [cs, stat].

\bibitem[{Sennrich et~al.(2016)Sennrich, Haddow, and Birch}]{bpe_paper}
Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016.
\newblock \href {https://doi.org/10.18653/v1/P16-1162} {Neural machine translation of rare words with subword units}.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, pages 1715--1725, Berlin, Germany. Association for Computational Linguistics.

\bibitem[{Shahshahani et~al.(2018)Shahshahani, Mohseni, Shakery, and Faili}]{peyma_ds}
Mahsa~Sadat Shahshahani, Mahdi Mohseni, Azadeh Shakery, and Heshaam Faili. 2018.
\newblock \href {https://arxiv.org/abs/1801.09936} {Peyma: A tagged corpus for persian named entities}.
\newblock \emph{Preprint}, arXiv:1801.09936.

\bibitem[{Sharami et~al.(2020)Sharami, Sarabestani, and Mirroshandel}]{dsp_ds}
Javad PourMostafa~Roshan Sharami, Parsa~Abbasi Sarabestani, and Seyed~Abolghasem Mirroshandel. 2020.
\newblock \href {https://arxiv.org/abs/2004.05328} {Deepsentipers: Novel deep learning models trained over proposed augmented persian sentiment corpus}.
\newblock \emph{Preprint}, arXiv:2004.05328.

\bibitem[{Touvron et~al.(2023)Touvron, Martin, Stone, Albert, Almahairi, Babaei, Bashlykov, Batra, Bhargava, Bhosale, Bikel, Blecher, Ferrer, Chen, Cucurull, Esiobu, Fernandes, Fu, Fu, Fuller, Gao, Goswami, Goyal, Hartshorn, Hosseini, Hou, Inan, Kardas, Kerkez, Khabsa, Kloumann, Korenev, Koura, Lachaux, Lavril, Lee, Liskovich, Lu, Mao, Martinet, Mihaylov, Mishra, Molybog, Nie, Poulton, Reizenstein, Rungta, Saladi, Schelten, Silva, Smith, Subramanian, Tan, Tang, Taylor, Williams, Kuan, Xu, Yan, Zarov, Zhang, Fan, Kambadur, Narang, Rodriguez, Stojnic, Edunov, and Scialom}]{llama_paper}
Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, Dan Bikel, Lukas Blecher, Cristian~Canton Ferrer, Moya Chen, Guillem Cucurull, David Esiobu, Jude Fernandes, Jeremy Fu, Wenyin Fu, Brian Fuller, Cynthia Gao, Vedanuj Goswami, Naman Goyal, Anthony Hartshorn, Saghar Hosseini, Rui Hou, Hakan Inan, Marcin Kardas, Viktor Kerkez, Madian Khabsa, Isabel Kloumann, Artem Korenev, Punit~Singh Koura, Marie-Anne Lachaux, Thibaut Lavril, Jenya Lee, Diana Liskovich, Yinghai Lu, Yuning Mao, Xavier Martinet, Todor Mihaylov, Pushkar Mishra, Igor Molybog, Yixin Nie, Andrew Poulton, Jeremy Reizenstein, Rashi Rungta, Kalyan Saladi, Alan Schelten, Ruan Silva, Eric~Michael Smith, Ranjan Subramanian, Xiaoqing~Ellen Tan, Binh Tang, Ross Taylor, Adina Williams, Jian~Xiang Kuan, Puxin Xu, Zheng Yan, Iliyan Zarov, Yuchen Zhang, Angela Fan, Melanie Kambadur, Sharan Narang, Aurelien Rodriguez, Robert Stojnic, Sergey Edunov, and Thomas
  Scialom. 2023.
\newblock \href {https://arxiv.org/abs/2307.09288} {Llama 2: Open foundation and fine-tuned chat models}.
\newblock \emph{Preprint}, arXiv:2307.09288.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin}]{vaswani_attention_2017}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan~N Gomez, Ł~ukasz Kaiser, and Illia Polosukhin. 2017.
\newblock \href {https://papers.nips.cc/paper_files/paper/2017/hash/3f5ee243547dee91fbd053c1c4a845aa-Abstract.html} {Attention is {All} you {Need}}.
\newblock In \emph{Advances in {Neural} {Information} {Processing} {Systems}}, volume~30. Curran Associates, Inc.

\bibitem[{Wang et~al.(2020)Wang, Pruksachatkun, Nangia, Singh, Michael, Hill, Levy, and Bowman}]{superglue_ds}
Alex Wang, Yada Pruksachatkun, Nikita Nangia, Amanpreet Singh, Julian Michael, Felix Hill, Omer Levy, and Samuel~R. Bowman. 2020.
\newblock \href {https://arxiv.org/abs/1905.00537} {Superglue: A stickier benchmark for general-purpose language understanding systems}.
\newblock \emph{Preprint}, arXiv:1905.00537.

\bibitem[{Wang et~al.(2019)Wang, Singh, Michael, Hill, Levy, and Bowman}]{glue_ds}
Alex Wang, Amanpreet Singh, Julian Michael, Felix Hill, Omer Levy, and Samuel~R. Bowman. 2019.
\newblock \href {https://openreview.net/forum?id=rJ4km2R5t7} {{GLUE}: A multi-task benchmark and analysis platform for natural language understanding}.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Zeng et~al.(2022)Zeng, Li, Wu, Liu, Liu, Yu, and Ma}]{unpading_paper}
Jinle Zeng, Min Li, Zhihua Wu, Jiaqi Liu, Yuang Liu, Dianhai Yu, and Yanjun Ma. 2022.
\newblock \href {https://arxiv.org/abs/2208.08124} {Boosting distributed training performance of the unpadded bert model}.
\newblock \emph{Preprint}, arXiv:2208.08124.

\bibitem[{Ziabary(2024)}]{targoman_pws}
S.~Mehran~M. Ziabary. 2024.
\newblock {T}argoman/{P}ersian{W}eb{S}craper: {A}n accurate scrapper to scrape popular persian websites, mostly intended to be used as a tool to create large corpora for {P}ersian language. --- github.com.
\newblock \url{https://github.com/Targoman/PersianWebScraper}.

\end{thebibliography}

\pdfoutput=1
\documentclass[11pt]{article}

\usepackage[preprint]{acl}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc}

\usepackage{microtype}

\newcommand\eztb[1]{\begin{tabular}[t]{@{}l@{}}#1\end{tabular}}

\usepackage{tabularray}

\usepackage{inconsolata}

\usepackage{graphicx}

\title{TookaBERT: A Step Forward for Persian NLU}

\author{
 \textbf{MohammadAli SadraeiJavaheri},
 \textbf{Ali Moghaddaszadeh},
 \textbf{Milad Molazadeh},
 \textbf{Fariba Naeiji},
 \\
 \textbf{Farnaz Aghababaloo},
 \textbf{Hamideh Rafiee},
 \textbf{Zahra Amirmahani},
 \textbf{Tohid Abedini},
 \\
 \textbf{Fatemeh Zahra Sheikhi},
 \textbf{Amirmohammad Salehoof}.
 \\
 Part Artificial Intelligence Research Center
}

\begin{document}
\maketitle
\begin{abstract}
The field of natural language processing (NLP) has seen remarkable advancements, thanks to the power of deep learning and foundation models. Language models, and specifically BERT, have been key players in this progress. In this study, we trained and introduced two new BERT models using Persian data. We put our models to the test, comparing them to seven existing models across 14 diverse Persian natural language understanding (NLU) tasks. The results speak for themselves: our larger model outperforms the competition, showing an average improvement of at least $+2.8$ points. This highlights the effectiveness and potential of our new BERT models for Persian NLU tasks.
\end{abstract}

\section{Introduction}

Recent advancements in natural language processing (NLP) are largely attributed to deep learning methods and the evolution of foundation models. Language models have played a pivotal role in this progress, with a standard approach of first training a language model and then fine-tuning it for specific NLP tasks\citep{howard-ruder-2018-universal, peters_deep_2018}. This method has become very popular in the field. Specifically after The transformer architecture, introduced by \citet{vaswani_attention_2017}, revolutionized these models. BERT, developed by \citet{devlin_bert_2019}, and GPT, created by \citet{Radford2018ImprovingLU}, were the first successful models to combine the transformer architecture with pre-training as a language model. A key difference between these two models is their approach to attention; GPT utilizes causal attention, while BERT employs a bidirectional approach. This distinction makes the GPT model more suitable for generative tasks.

Generative tasks quickly embraced GPT as the model of choice, while understanding tasks found greater success with BERT. The release of GPT-3 \citep{NEURIPS2020_1457c0d6} further boosted the popularity of generative models in natural language processing (NLP). With the introduction of ChatGPT, generative models have taken center stage. These models require a large number of parameters to perform well across different tasks. While they offer impressive capabilities and versatility, they often necessitate substantial computational power and resources. Training or fine-tuning such large models on standard computers is not feasible.

In contrast, smaller encoder models have demonstrated remarkable performance in natural language understanding tasks. These models can be fine-tuned and executed with far less hardware, and their performance after fine-tuning is comparable to that of a large language model on specific tasks.

In this study, we address the lack of a large-scale Persian BERT model by training a new model using state-of-the-art techniques on a large Persian dataset. We also trained a base version to serve as a comparison point. Our main contributions are twofold: \textbf{(I)} We are making available to the public two new checkpoints of the BERT model specifically trained for the Persian language, which fills a significant gap in the availability of large-scale models for this language; \textbf{(II)} Through systematic evaluation across 14 diverse Persian natural language understanding (NLU) tasks, our larger checkpoint demonstrates superior performance, achieving an average improvement of at least $+2.8$ points compared to other models.

\section{Related Works}

\begin{table*}[t]

\begin{center}
\resizebox{\textwidth}{!}{

\begin{tblr}{
    column{2-9} = {c},
    vline{2, 7} = {1-3}{},
    vline{9} = {-}{},
    cell{2}{2,3,4} = {}{b},
    cell{1}{2} = {c=5}{c},
    cell{1}{7} = {c=2}{c},
    hline{2, 11} = {-}{},
    hline{4, 13} = {-}{2px},
}
  & ParsiNLU &&&&& NER \\
Task→ & \eztb{Reading\\Comprehension} & \eztb{Sentiment\\Analysis} & QQP & \eztb{Multiple\\Choice} & \eztb{Entailment} & \eztb{Peyma NER} & \eztb{MultiConer} & Avg.\\
Models↓  Metrics→ & EM/F1 & F1/Acc. & F1/Acc. & F1/Acc. & F1/Acc. & F1/Acc. & F1/Acc. & -\\
XLM-V & 8.0/26.7 & 82.3/84.1 & 80.1/80.0 & 35.3/35.2 & 28.3/37.8 & \textbf{88.1/97.8} & 58.8/92.2 & 59.6\\
XLM-RoBERTa-Base & 20.0/40.4 & 89.7/89.9 & 79.1/79.0 & 32.4/32.4 & 53.8/53.7 & 87.8/97.8 & 60.4/92.5 & 64.9\\
Bert Multilingual & 27.2/42.2 & 86.8/87.5 & 79.2/79.3 & 33.4/33.1 & 54.4/54.4 & 83.7/96.9 & 60.3/92.5 & 65.1\\
Shiraz & 17.6/39.6 & 87.1/87.8 & 79.7/79.5 & 34.7/34.5 & 39.9/41.5 & 86.4/97.6 & 59.1/92.8 & 62.7\\
Parsbert & 20.0/39.6 & 88.6/88.9 & 80.2/80.1 & 35.3/35.2 & 49.9/49.6 & 86.8/97.7 & 64.9/93.2 & 65.0\\
AriaBERT & 14.4/35.5 & 87.9/88.4 & 79.1/78.8 & 30.8/30.9 & 44.5/44.4 & 84.1/97.3 & 61.0/92.5 & 62.1\\
FaBERT & 27.2/48.4 & 90.4/90.5 & 82.3/82.3 & 32.5/32.4 & \textbf{55.0/54.8} & 87.4/97.6 & 63.9/93.0 & 67.0\\
TookaBERT-Base & 20.8/42.5 & 89.6/89.7 & 81.3/81.3 & 33.6/33.8 & 49.3/48.9 & 86.3/97.8 & 66.2/93.3 & 65.3\\
TookaBERT-Large & \textbf{33.6/60.5} & \textbf{91.1/91.4} & \textbf{82.7/82.6} & \textbf{36.1/36.0} & 54.3/54.1 & 86.2/98.0 & \textbf{69.7/94.1} & \textbf{69.3}\\

\end{tblr}

}
\end{center}

\caption{\label{table_r1}
Performance comparison of various base models on different end-tasks. The 'Average' column is calculated based on the scores presented in this table for both ParsiNLU and named entity recognition tasks.}
\end{table*}

BERT was introduced by \citep{devlin_bert_2019} and improved by RoBERTa \citep{liu_roberta_2019}. The key difference between these models is their training objective. RoBERTa removed the next sentence prediction loss, focusing solely on masked language modeling loss, and demonstrated that this alone is sufficient for training.

There are several pre-trained Persian BERT models, with the earliest and most well-known being ParsBERT by \citet{parsbert_model}. Other models include AriaBERT\citep{ariabert_model}, FaBERT\citep{masumi_fabert_2024}, and the recent Shiraz and Tehran models by \citet{shiraz_tehran}. While Persian BERT models from last year outperformed ParsBERT, none of them experimented with model sizes beyond BERT-Base, such as BERT-Large. Multilingual pre-trained BERT models have also exhibited good performance on Persian tasks \citep{conneau_unsupervised_2020, liang_xlm-v_2023}.

\section{Data}

\noindent\textbf{Pre-training dataset:} This includes a mix of public datasets: hmblogs \citep{khansari2021hmblogs}, the cleaned Persian subset of madlad \citep{kudugunta2023madlad400}, and PersianWebScrapper \citep{targoman_pws}. We normalized these datasets using NFKC to ensure standard UTF-8 encoding for Persian text. We also replaced some Arabic characters with their Persian equivalents.

\noindent\textbf{Evaluation dataset:} We aggregated multiple Persian NLU datasets for evaluation. Unlike English, which has well-known collections like GLUE \citep{glue_ds} and SuperGLUE \citep{superglue_ds}, Persian lacks a comprehensive NLU dataset collection. We used the ParsiNLU dataset \citep{parsinlu_ds}, which includes 6 tasks, one of which is machine translation and is not relevant for BERT evaluation. Therefore, we utilized the remaining 5 tasks from ParsiNLU along with the following individual datasets: PeymaNER \citep{peyma_ds}, Persian subset of MultiCoNER \citep{multiconer_ds}, PQA \citep{pqa_ds}, PQuAD \citep{pquad_ds}, FarsTail \citep{farstail_ds}, DeepSentiPers \citep{dsp_ds}, SnappFood sentiment dataset\footnote{Available at: \url{https://bit.ly/2Xu2xq1}} \citep{parsbert_model}, SentiPers \citep{sp_ds}, and Arman Emotion \citep{arman_emotion}. Our aim is to give a fair comparison for the foundation models we are testing by using a variety of Persian language tasks.

\section{Methodology}

\begin{table*}[t]

\begin{center}
\resizebox{\textwidth}{!}{

\begin{tblr}{
    column{2-9} = {c},
    cell{1}{2} = {c=2}{c},
    cell{1}{4} = {}{c},
    cell{1}{5} = {c=4}{c},
    vline{2, 4, 5} = {1-3}{},
    vline{9} = {-}{},
    hline{2} = {2-9}{},
    hline{2, 11} = {-}{},
    hline{4, 13} = {-}{2px},
}
  & QA &&NLI & Sentiment &&\\
Task→ & \eztb{PQA} & \eztb{PQuAD} & \eztb{FarsTail} & \eztb{DeepSentiPers} & \eztb{Snappfood} & \eztb{SentiPers} & \eztb{Arman\\Emotion} & Avg.\\
Models↓  Metrics→ & EM/F1/Has-EM/Has-F1 & EM/F1/Has-EM/Has-F1 & F1/Acc. & F1/Acc. & F1/Acc. & F1/Acc. & F1/Acc. & -\\
XLM-V & 40.3/57.0/29.8/62.3 & 73.3/85.7/68.2/85.6 & 81.1/81.2 & 83.4/83.4 & 88.0/88.1 & 74.6/74.8 & 42.3/48.0 & 71.1\\
XLM-RoBERTa-Base & 39.2/58.0/30.2/62.5 & 73.7/86.2/68.2/85.8 & 82.0/82.0 & 84.0/84.1 & 87.6/87.7 & 75.7/75.8 & 69.8/70.0 & 75.0\\
Bert Multilingual & 37.8/56.0/28.1/60.9 & 71.8/84.7/65.9/84.0 & 82.7/82.8 & 78.6/78.7 & 87.1/87.1 & 71.7/71.9 & 62.2/62.4 & 72.1\\
Shiraz & 30.4/44.4/18.1/51.6 & 66.0/81.2/59.6/81.3 & 77.8/77.8 & 81.2/81.1 & 87.1/87.1 & 73.4/73.8 & 71.6/71.7 & 71.3\\
Parsbert & 36.7/49.8/26.8/56.1 & 71.4/84.2/66.3/84.6 & 80.9/80.9 & 80.2/80.2 & 87.6/87.7 & 74.3/74.4 & 67.2/67.4 & 72.8\\
AriaBERT & 35.0/44.4/19.5/48.1 & 68.1/81.2/62.1/80.9 & 74.5/74.4 & 80.5/80.5 & 87.7/87.8 & 73.6/73.8 & 71.4/71.5 & 71.1\\
FaBERT & 39.8/55.6/30.0/62.2 & 72.6/85.4/67.2/85.3 & 83.7/83.7 & 82.7/82.7 & \textbf{88.1/88.1} & 75.0/75.3 & 72.9/73.0 & 75.3\\
TookaBERT-Base & 38.7/54.4/29.2/60.9 & 73.2/85.7/68.3/85.9 & 83.3/83.4 & 83.9/83.9 & 87.4/87.4 & 74.9/75.1 & 73.2/73.4 & 75.3\\
TookaBERT-Large & \textbf{46.2/65.0/38.4/69.9} & \textbf{75.6/88.1/70.2/87.8} & \textbf{89.7/89.7} & \textbf{85.7/85.8} & \textbf{88.1/88.1} & \textbf{76.3/76.5} & \textbf{74.7/74.7} & \textbf{78.6}\\

\end{tblr}

}
\end{center}

\caption{\label{table_r2}
Performance comparison of base models on various end-tasks. The 'Average' column is calculated based on the scores from question answering, natural language inference, and sentiment analysis datasets, as presented in this table.}
\end{table*}

\subsection{Tokenizer}
The original BERT paper used the WordPiece algorithm for tokenization \citep{devlin_bert_2019}. However, newer models mostly use byte pair encoding (BPE) \citep{bpe_paper}. Examples include RoBERTa \citep{liu_roberta_2019} and LLaMA \citep{llama_paper}. We also chose to use BPE for our tokenizer. We trained a new Persian tokenizer using this algorithm and set our vocab size to 48,000. During the training phase, we also used BPE-dropout \citep{bpe-drop} of 0.1 to regularize our model and improve generalization over unseen tokens.

\subsection{BERT Pre-training}
We trained our BERT model using the latest advancements and ideas in the field. Our approach was influenced by other works such as MosaicBERT \citep{portes2023mosaicbert}. To speed up training and efficiently utilize GPU memory, we employed flash attention v2 \citep{dao_flashattention-2_2023}. Additionally, we made use of multiple GPUs effectively by utilizing the ZeRO stage 2 optimizer \citep{rajbhandari_zero_2020}.

Similar to RoBERTa \citep{liu_roberta_2019}, we used only mask language modeling as our objective and excluded next sentence prediction. To avoid the computational waste of padding tokens, we concatenated our entire dataset and used it as a single batch \citep{unpading_paper}. We also incorporated whole-word-masking \citep{cui_pre-training_2021} to make predicting masks more challenging and improve the model's capabilities.

We trained two BERT models in standard \textbf{Base} and \textbf{Large} sizes. The training was conducted using 8xA100 40GB GPUs.

\subsection{Evaluation}
For a fair comparison, we fine-tuned each model with four different learning rates: 3e-5, 3e-6, 7e-5, and 7e-6 for eight epochs. We then reported the best evaluation score achieved by each model across these learning rates and epochs. In addition to our models, we also evaluated the following pre-trained language models: ParsBERT \citep{parsbert_model}, AriaBERT \citep{ariabert_model}, FaBERT \citep{masumi_fabert_2024}, Shiraz \citep{shiraz_tehran}, BERT multilingual \citep{devlin_bert_2019}, XLM-RoBERTa-Base \citep{conneau_unsupervised_2020}, and XLM-V \citep{liang_xlm-v_2023}

\section{Results}
We employed various evaluation metrics depending on the nature of the task. The score for each task was computed by averaging the respective metrics, with the F1-score and accuracy being the most prevalent across all tasks.

The results, as shown in Tables \ref{table_r1} and  \ref{table_r2}, indicate that TookaBERT-Large demonstrated superior performance compared to other models in the majority of end-tasks. On average, it achieved an impressive 2.8-point improvement over models such as FaBERT.

\section{Conclusion}

In this study, we presented two new BERT models specifically trained for Persian natural language understanding (NLU) tasks. By employing state-of-the-art techniques and a large Persian dataset, we successfully trained and introduced TookaBERT-Base and TookaBERT-Large to address the lack of large-scale Persian BERT models. Our comprehensive evaluations across 14 diverse Persian NLU tasks showcased the superiority of our larger model, outperforming other existing models by an average of $+2.8$ points. This highlights the effectiveness and potential of our new BERT models for Persian language processing.

The significant performance of TookaBERT-Large underscores the importance of model size and training techniques in language model development. By making our models publicly available, we hope to contribute to the advancement of Persian NLU and encourage further research in this direction. Our work showcases the potential of adapting foundation models to specific languages, and we believe that future studies can build upon our models to further enhance their performance and applicability in various Persian language tasks.

\bibliography{custom}

\end{document}